### v2

@@ Abstract
Integration of unmanned aerial vehicles (UAVs) for surveillance or monitoring applications into fifth generation (5G) New Radio (NR) cellular networks is an intriguing problem that has recently tackled a lot of interest in both academia and industry. For an efficient spectrum usage, we consider a recently-proposed sky-ground nonorthogonal multiple access (NOMA) scheme, where a cellular-connected UAV acting as aerial user (AU) and a static terrestrial user (TU) are paired to simultaneously transmit their uplink signals to a base station (BS) in the same time-frequency resource blocks. In such a case, due to the highly dynamic nature of the UAV, the signal transmitted by the AU experiences both time dispersion due to multipath propagation effects and frequency dispersion caused by Doppler shifts. On the other hand, for a static ground network, frequency dispersion of the signal transmitted by the TU is negligible and only multipath effects have to be taken into account. To decode the superposed signals at the BS through successive interference cancellation, accurate estimates of both the AU and TU channels are needed. In this paper, we propose channel estimation procedures that suitably exploit the different circular/noncircular modulation formats (modulation diversity) and the different almost-cyclostationarity features (Doppler diversity) of the AU and TU by means of widely-linear time-varying processing. Our estimation approach is semi-blind since Doppler shifts and time delays of the AU are estimated based on the received data only, whereas the remaining relevant parameters of the AU and TU channels are acquired relying also on the available training symbols. Monte Carlo numerical results demonstrate that the proposed channel estimation algorithms can satisfactorily acquire all the relevant parameters in different operative conditions.
@@
Almost-cyclostationarity, channel estimation, circular and noncircular modulations, double-selective channels, Doppler diversity, 5G New Radio (NR), modulation diversity, non-orthogonal multiple access (NOMA), unmanned aerial vehicles (UAVs), uplink.
@@ 1 Introduction
With the development of unmanned aerial vehicles (UAVs), computer vision, and sensor technology, UAV systems have been increasingly employing in civilian and commercial applications, such as surveillance and monitoring, due to their ability to quickly cover large and difficult-to-reach areas. The recently-introduced concept of cellular-connected UAVs [1, 2] represents a viable strategy to integrate UAVs in fifth generation (5G) New Radio (NR) networks, by allowing them to transmit data to the cellular network as aerial users (AUs). 3GPP Release 18 is introducing 5G NR support to enhance the safe use of UAVs for commercial and leisure applications. Nevertheless, the integration of UAVs into cellular networks introduces two interelated basic challenges [3, 4]. First, existing 5G cellular infrastructures are designed to serve terrestrial users (TUs) and, thus, the antennas of the base stations (BSs) look downward: such a tilting might cause AUs to be served by the side lobes of the BSs and, consequently, they suffer from a reduction in antenna gain. Second, due to their mobility capabilities, AUs have more versatile movements than TUs, thereby introducing Doppler (frequency) shifts, which result in carrier frequency offset, inter-carrier interference, and reduced channel coherence time.
@@ 1.1 NOMA for cellular-connected UAVs
The works [32, 33] deserve a separate discussion. Orthogonal time frequency space (OTFS) modulation has been used in [32] for the implementation of NOMA among users with different mobility profiles: specifically, a high-mobility user is allowed to share the spectrum with multiple low-mobility NOMA users. Although the specific case of an AU is not explicitly considered in [32, 33], their framework may be adapted to the scenario of cellular-connected UAVs, by regarding the high-mobility user as an UAV. However, besides entailing more complex signal processing techniques at both the transmitter and receiver, compared to standard orthogonal frequency-division multiplexing (OFDM) modems, zero fractional delays and zero fractional Doppler shifts are assumed in the OTFS-NOMA scheme. Last but not least, Doppler shifts are still seen as a source of performance degradation to be counteracted and perfect channel knowledge is assumed in [32, 33].
Channel estimation specifically targeted at cellular-connected UAVs remains a critical issue, and effective methods operating over shared aerial and terrestrial channels are an open research issue.
@@ 1.2 Contribution and organization
We consider a NOMA scheme, which is referred to as sky-ground (SG) NOMA, where a flying AU is paired with a static (i.e., fixed) TU. The AU and TU have different propagation conditions and their received power can be rather different at the BS, regardless of their transmit power: indeed, the air-to-cellular (A2C) channel between the AU and the elevated terrestrial BS is both time- and frequency-selective (i.e., doubly-selective) with significant Doppler shifts, [^1] The air-to-cellular channel is different from the air-to-ground one due to the non-negligible height of terrestrial BSs. whereas the TU-to-BS channel is predominately frequency-selective with negligible Doppler effects. We propose to exploit the different mobile profiles of the AU and TU rather than to counteract the Doppler effects induced by the mobility of the UAV. In particular, the AU is allowed to continuously transmit along all its trajectory.
@@
Acquisition of channel state information (CSI) in uplink NOMA is a challenging task, since channel estimation in current wireless data communications standards is mainly based on training sequences sent from the users to the BS that are orthogonal in either time or frequency domain. Orthogonal RBs for training adversely affect spectral efficiency and system performance, especially in the uplink. We consider a training scheme where the training data of one user is contaminated by either the pilots or the information-bearing data of the other user. In this challenging scenario, we propose a two-stage approach: first, the channel parameters of the AU are semi-blindly estimated by also exploiting the almost-cyclostationarity (ACS) properties [34] of the received signal that depend on the Doppler shifts of the AU-to-BS channel (Doppler diversity), as well as the different circular/noncircular features of the AU and TU (modulation diversity); subsequently, capitalizing on the CSI knowledge of the AU, the overall channel of the TU is acquired with a pilot-aided improved widely-linear (WL) estimator [35]. The performances of the proposed channel estimation algorithms are validated through extensive numerical examples.
The paper is organized as follows. The system model of the considered uplink SG-NOMA is described in Section 2. The proposed estimation strategies for the AU and TU are developed in Sections 3 and 4, respectively. Numerical results are reported in Section 5 in terms of channel estimation accuracy. Finally, the main results obtained in the paper are summarized in Section 6.
@@ 1.3 Basic notations
Upper- and lower-case bold letters denote matrices and vectors; the superscripts ***, T, H, and $-1$ denote the conjugate, the transpose, the Hermitian (conjugate transpose), and the inverse of a matrix; $\mathbb{C}$, $\mathbb{R}$ and $\mathbb{Z}$ are the fields of complex, real and integer numbers; $\mathbb{C}^{n}$ $[\mathbb{R}^{n}]$ denotes the vector-space of all $n$-column vectors with complex [real] coordinates; similarly, $\mathbb{C}^{n\times m}$ $[\mathbb{R}^{n\times m}]$ denotes the vector-space of all the $n\times m$ matrices with complex [real] elements; $\delta(\tau)$ is the Dirac delta; $\delta_{n}$ is the Kronecker delta, i.e., $\delta_{n}=1$ when $n=0$ and zero otherwise; $\jmath\triangleq\sqrt{-1}$ denotes the imaginary unit; $\text{max}(x,y)$ [$\text{min}(x,y)$] returns the maximum [minimum] between $x\in\mathbb{R}$ and $y\in\mathbb{R}$; $\lfloor x\rfloor$ rounds $x\in\mathbb{R}$ to the nearest integer less than or equal to $x$; the (linear) convolution operator is denoted with $\ast$; $\frac{\partial}{\partial x*}f$ is the complex derivative of the real-valued scalar function $f$ with respect to $x^{*}$ [36]; $\mathbf{0}_{n}$, ${\mathbf{O}}_{n\times m}$ and ${\mathbf{I}}_{n}$ denote the $n$-column zero vector, the $n\times m$ zero matrix and the $n\times n$ identity matrix; ${\mathbf{J}}_{n}$ is given by
@@ 2 System model and basic assumptions
The considered single-cell network scenario encompasses an elevated BS, mobile AUs, and static TUs. The BS is located within the cell at a fixed height $H_{\text{BS}}$, whereas the AU flies at an altitude $H_{\text{A}}$ above the ground. [^2] The proposed estimation algorithms can be straightforwardly modified to account for the case where the AU varies its height along the trajectory. We restrict ourselves to uplink communications since advanced signal detection algorithms of NOMA are more affordable at the BS [24]. A uniform linear array (ULA) is used at the BS with $J$ antennas. The AUs are represented by single-antenna UAVs that are employed for surveillance or monitoring a given area within a cell and transmit data to the corresponding BS over given A2C links. According to existing channel access schemes for terrestrial cellular networks, we assume that TUs transmit to the BS using the OMA scheme and, hence, there is no interference among the TUs in the cell.
In this paper, we consider an uplink SG-NOMA protocol, where single-antenna TUs share their RBs with the AUs. As in [30], we assume that each AU is randomly paired with a different static (i.e., fixed) TU (see Fig. 1). The two-user scenario represents the case of major practical interest for NOMA, since it results in acceptable interference levels [38].
@@ 2.2 Channel model for terrestrial user
${\mathbf{a}}(\theta)\triangleq\left(1,e^{\jmath\frac{2\pi}{\lambda_{\text{carrier}}}d\sin(\theta)},\ldots,e^{\jmath\frac{2\pi}{\lambda_{\text{carrier}}}(J-1)d\sin(\theta)}\right)^{\text{T}}\in\mathbb{C}^{J}$ (5)
@@ 2.3 Channel model for the aerial user
For A2C link, we assume a two-ray channel model, which includes the LoS path and a scattered (i.e., non-LoS) component [8] (see also Fig. 1). The non-LoS (NLoS) component is physically due to many rays reflected or scattered from closely-spaced terrestrial obstacles within the vicinity of the BS, which are characterized by a very narrow beamwidth [8] and appear grouped or “clustered" in delay for sufficiently high elevation angles [44]. Henceforth, such rays are expected to have a very similar Doppler shift, as they all come from similar directions (narrow beamwidth), and a very similar delay (high elevation angle). Assuming that the UAV is moving (relative to the BS) at constant radial speed $v$ within the observation interval $[0,T_{0})$, with $T_{0}\triangleq N_{0}\,T_{\text{s }}\leq T_{\text{coh}}$, under the customary assumption that the communication bandwidth $B\sim 1/T_{\text{c}}$ is much smaller than the carrier frequency $f_{\text{carrier}}$, the far-field low-pass equivalent linear time-varying (LTV) time-delay impulse response of the UAV-to-BS channel can be expressed as [^4] Throughout the paper, the subscripts L and N indicate parameters referring to the LoS and NLoS components of the AU, respectively. (see, e.g., [45])
@@ 2.4 Receive signal model before CP removal
$x_{\text{A,C}}^{[p]}[n]$ $=\sum_{\ell=n-1}^{n}\sum_{q=0}^{P-1}u_{\text{A}}^{[q]}[\ell]\,\alpha_{\text{A,C}}\left[(n-\ell)P+(p-q)-d_{\text{A,C}}\right]$ (11)
$x_{\text{T},k}^{[p]}[n]$ $=\sum_{\ell=n-1}^{n}\sum_{q=0}^{P-1}u_{\text{T}}^{[q]}[\ell]\,\alpha_{\text{T},k}\left[(n-\ell)P+(p-q)-d_{\text{T},k}\right]$ (12)

### v4

@@ Abstract
Integration of unmanned aerial vehicles (UAVs) for surveillance or monitoring applications into fifth generation (5G) New Radio (NR) cellular networks is an intriguing problem that has recently tackled a lot of interest in both academia and industry. For an efficient spectrum usage, we consider a recently-proposed sky-ground nonorthogonal multiple access (NOMA) scheme, where a cellular-connected UAV acting as aerial user (AU) and a static terrestrial user (TU) are paired to simultaneously transmit their uplink signals to a base station (BS) in the same time-frequency resource blocks. In such a case, due to the highly dynamic nature of the UAV, the signal transmitted by the AU experiences both time dispersion due to multipath propagation effects and frequency dispersion caused by Doppler shifts. On the other hand, for a static ground network, frequency dispersion of the signal transmitted by the TU is negligible and only multipath effects have to be taken into account. To decode the superposed signals at the BS through successive interference cancellation, accurate estimates of both the AU and TU channels are needed. In this paper, we propose channel estimation procedures that suitably exploit the different circular/noncircular modulation formats (modulation diversity) and the different almost-cyclostationarity features (Doppler diversity) of the AU and TU by means of widely-linear time-varying processing. Our estimation approach is semi-blind since Doppler shifts and time delays of the AU are estimated based on the received data only, whereas the remaining relevant parameters of the AU and TU channels are acquired relying also on the available training symbols, which are transmitted by the AU and TU in a nonorthogonal manner. Monte Carlo numerical results demonstrate that the proposed channel estimation algorithms can satisfactorily acquire all the relevant parameters in different operative conditions.
@@
Almost-cyclostationarity, channel estimation, circular and noncircular modulations, doubly selective channels, Doppler diversity, 5G New Radio (NR), modulation diversity, non-orthogonal multiple access (NOMA), unmanned aerial vehicles (UAVs), uplink.
This article has been accepted for publication in a future issue of this journal, but has not been fully edited. Content may change prior to final publication.
Citation information: DOI 10.1109/OJCOMS.2024.3451308, IEEE Open Journal of Communications Society.
@@ 1 Introduction
With the development of unmanned aerial vehicles (UAVs), computer vision, and sensor technology, UAV systems have been increasingly employing in civilian and commercial applications, such as surveillance and monitoring, due to their ability to quickly cover large and difficult-to-reach areas. The recently-introduced concept of cellular-connected UAVs [1, 2, 3] represents a viable strategy to integrate UAVs in fifth generation (5G) New Radio (NR) networks, by allowing them to transmit data to the cellular network as aerial users (AUs). 3GPP Release 18 is introducing 5G NR support to enhance the safe use of UAVs for commercial and leisure applications. Nevertheless, the integration of UAVs into cellular networks introduces two interelated basic challenges [4, 5]. First, existing 5G cellular infrastructures are designed to serve terrestrial users (TUs) and, thus, the antennas of the base stations (BSs) look downward: such a tilting might cause AUs to be served by the side lobes of the BSs and, consequently, they suffer from a reduction in antenna gain. Second, due to their mobility capabilities, AUs have more versatile movements than TUs, thereby introducing Doppler (frequency) shifts, which result in carrier frequency offset, inter-carrier interference, and reduced channel coherence time.
@@ 1.1 NOMA for cellular-connected UAVs
1.2 Modulation formats for double-selective channels
The choice of the digital modulation format crucially determines accuracy and signaling overhead of channel estimation. Orthogonal time frequency space (OTFS) and orthogonal frequency division multiplexing (OFDM) are digital modulation formats designed to handle doubly (i.e., time-and-frequency) selective channels. OFDM modulation directly places symbols in the time-frequency (TF) domain, whereas OTFS is a modulation technique in which symbols are transmitted in the delay-Doppler (DD) domain [33].
OFDM is immune to the frequency-selective nature of the wireless channel. However, if frequency dispersion is also introduced by the channel in the form of Doppler shifts, the orthogonality of the subcarriers is destroyed, thus resulting in intercarrier interference (ICI). On the other hand, OTFS might cope with multipath fading and Doppler shifts more effectively, but such a robustness comes at the cost of a nonnegligible increase in signal processing complexity.
OTFS modulation has been used in [34] for the implementation of NOMA among users with different mobility profiles: specifically, a high-mobility user is allowed to share the spectrum with multiple low-mobility NOMA users. Although the specific case of an AU is not explicitly considered in [34, 35], their framework may be adapted to the scenario of cellular-connected UAVs, by regarding the high-mobility user as an UAV. However, perfect channel knowledge is assumed in [34, 35].
The accuracy and signaling overhead of channel estimation techniques for OTFS modulation strictly depend on the values assumed by the delays and Doppler shifts. Indeed, in the case of integer delays and Doppler shifts, the received symbol corresponding to a given channel path is well-localized in the corresponding DD bin [34, 35, 36]. On the other hand, for fractional delays and Doppler shifts, each transmitted symbol spreads into adjacent bins, resulting in received symbols interfering with each other [37]. Such an inter-path interference (IPI) is a source of performance degradation for channel estimation performed directly in the DD domain. In the case of training-based channel identification, which consists into placing pilot symbols in the DD grid, one way to reduce the detrimental effects of IPI is to introduce guard symbols [38]. However, the guard interval of the pilots needs to include the entire Doppler domain within the range of the maximum delay, which results in a large pilot overhead. To avoid waste of communication resources, cancellation procedures of IPI might be carried out before channel estimation [37, 39], which inevitably increase implementation complexity of OTFS coherent reception.
Channel estimation for OFDM modulation is directly performed in the TF domain and, with respect to OTFS, it can face with fractional delays and Doppler shifts more effectively. Moreover, OFDM was extensively used in many standards of wireless systems, e.g., digital audio/video broadcasting (DAB/DVB), Wi-Fi local networks, worldwide interoperability for microwave access (WiMAX), long term evolution (LTE) and LTE-Advanced, and, it has been adopted in the 5G NR network. OFDM is still considered one of the potential candidates for beyond-5G communication systems. However, since OFDM is more sensitive to Doppler shifts than OTFS, a higher level of estimation accuracy is demanded. Despite the fact that channel estimation for OFDM is a well-investigated topic, to the best of our knowledge, there are no effective channel estimation methods available yet for NOMA aerial and terrestrial OFDM uplink channels.
@@ 1.3 Contribution and organization
We consider a NOMA uplink scheme, which is referred to as sky-ground (SG) NOMA, where a flying AU is paired with a static (i.e., fixed) TU. The AU and TU have different propagation conditions and their received power can be rather different at the BS, regardless of their transmit power: indeed, the air-to-cellular (A2C) channel between the AU and the elevated terrestrial BS is doubly selective with significant Doppler shifts, [^1] The air-to-cellular channel is different from the air-to-ground one due to the non-negligible height of terrestrial BSs. whereas the TU-to-BS channel is predominately frequency-selective with negligible Doppler effects. We propose to exploit the different mobile profiles of the AU and TU rather than to counteract the Doppler effects induced by the mobility of the UAV. In particular, the AU is allowed to continuously transmit along all its trajectory.
@@
Acquisition of channel state information (CSI) in uplink NOMA is a challenging task, since channel estimation in current wireless data communications standards is mainly based on training sequences sent from the users to the BS that are orthogonal in either time or frequency domain. Orthogonal RBs for training adversely affect spectral efficiency and system performance, especially in the uplink. We consider a nonorthogonal training scheme where the training data of one user is contaminated by either the pilots or the information-bearing data of the other user. In this challenging scenario, we propose a two-stage approach: first, the channel parameters of the AU are semi-blindly estimated by also exploiting the almost-cyclostationarity (ACS) properties [40] of the received signal that depend on the Doppler shifts of the AU-to-BS channel (Doppler diversity), as well as the different circular/noncircular features of the AU and TU (modulation diversity); subsequently, capitalizing on the CSI knowledge of the AU, the overall channel of the TU is acquired with a pilot-aided improved widely-linear (WL) estimator [41]. The performances of the proposed channel estimation algorithms are validated through extensive Monte Carlo numerical examples.
The paper is organized as follows. The system model of the considered uplink SG-NOMA is described in Section 2. The proposed estimation strategies for the AU and TU are developed in Sections 3 and 4, respectively. Monte Carlo numerical results are reported in Section 5 in terms of channel estimation accuracy. Finally, the main results obtained in the paper are summarized in Section 6.
@@ 1.4 Basic notations
Upper- and lower-case bold letters denote matrices and vectors; the superscripts *, T, H, and $-1$ denote the conjugate, the transpose, the Hermitian (conjugate transpose), and the inverse of a matrix; $\mathbb{C}$, $\mathbb{R}$ and $\mathbb{Z}$ are the fields of complex, real and integer numbers; $\mathbb{C}^{n}$ $[\mathbb{R}^{n}]$ denotes the vector-space of all $n$-column vectors with complex [real] coordinates; similarly, $\mathbb{C}^{n\times m}$ $[\mathbb{R}^{n\times m}]$ denotes the vector-space of all the $n\times m$ matrices with complex [real] elements; $\delta(\tau)$ is the Dirac delta; $\delta_{n}$ is the Kronecker delta, i.e., $\delta_{n}=1$ when $n=0$ and zero otherwise; $\jmath\triangleq\sqrt{-1}$ denotes the imaginary unit; $\text{max}(x,y)$ [$\text{min}(x,y)$] returns the maximum [minimum] between $x\in\mathbb{R}$ and $y\in\mathbb{R}$; $\lfloor x\rfloor$ rounds $x\in\mathbb{R}$ to the nearest integer less than or equal to $x$; the (linear) convolution operator is denoted with $\ast$; $\frac{\partial}{\partial x*}f$ is the complex derivative of the real-valued scalar function $f$ with respect to $x^{*}$ [42]; $\mathbf{0}_{n}$, ${\mathbf{O}}_{n\times m}$ and ${\mathbf{I}}_{n}$ denote the $n$-column zero vector, the $n\times m$ zero matrix and the $n\times n$ identity matrix; ${\mathbf{J}}_{n}$ is given by
@@ 2 System model and basic assumptions
The considered single-cell network scenario encompasses an elevated BS, mobile AUs, and static TUs. The BS is located within the cell at a fixed height $H_{\text{BS}}$, whereas the AU flies at an altitude $H_{\text{A}}$ above the ground. [^2] The proposed estimation algorithms can be straightforwardly modified to account for the case where the AU varies its height along the trajectory. We restrict ourselves to uplink communications since advanced signal detection algorithms of NOMA are more affordable at the BS [25]. A uniform linear array (ULA) is used at the BS with $J$ antennas. The AUs are represented by single-antenna UAVs that are employed for surveillance or monitoring a given area within a cell and transmit data to the corresponding BS over given A2C links. According to 5G NR channel access schemes for terrestrial cellular networks [44], we assume that TUs transmit to the BS using the OMA scheme and, hence, there is no interference among the TUs in the cell. The extension of the proposed approach to the case in which nonorthogonal RBs are assigned to TUs in the uplink is discussed in Subsections III-3.4 and IV-4.2.
In this paper, we consider an uplink SG-NOMA protocol, where single-antenna TUs share their RBs with the AUs. As in [31], we assume that each AU is randomly paired with a different static (i.e., fixed) TU (see Fig. 1). Hence, the scenario of interest boils down to the case of a single AU and a paired TU. [^3] In principle, multiple AUs can be paired with a single static TU. In this case, the proposed estimation framework might be extended straightforwardly, provided that the superimposed AU signals exhibit different ACS properties at the BS, i.e., the AU-to-BS channels are characterized by different Doppler shifts. The two-user scenario represents the case of major practical interest for NOMA, since it results in acceptable interference levels [45]. Moreover, we assume that inter-cell interference (ICI) coordination techniques are employed to mitigate terrestrial ICI [46, 47]. When the UAV flies in the low-altitude airspace [6, 7, 8], the LoS probability between the UAV and the neighboring co-channel BSs may be very low and ICI coordination (ICIC) only needs to involve a few BSs in a local region to exchange control information via high-speed backhaul links. In this case, the use of more sophisticated aerial-ground interference mitigation/cancellation procedures [3] might be avoided. Additionally, since each AU uses NOMA with a paired TU in the same cell, terrestrial ICIC indirectly limits aerial-ground interference between the UAV and co-channel terrestrial communications. Thus, in this work, we focus on the equivalent scenario with a single BS and a single AU-TU pairing in a given cell, without ICI.
@@ 2.2 Channel model for terrestrial user
${\mathbf{a}}(\theta)\triangleq\left(1,e^{\jmath\frac{2\pi}{\lambda_{\text{carrier}}}d\sin(\theta)},\ldots,e^{\jmath\frac{2\pi}{\lambda_{\text{carrier}}}(J-1)d\sin(\theta)}\right)^{\text{T}}$ (5)
@@ 2.3 Channel model for the aerial user
For A2C link, we assume a two-ray channel model, [^5] The proposed estimation framework can be extended to the multi-ray case for the A2C channel. However, for blind identification of the corresponding channel parameters, the Doppler shifts of the different rays have to fulfill a mild technical condition to avoid ambiguities [53]. which includes the LoS path and a scattered (i.e., non-LoS) component [9] (see also Fig. 1). The non-LoS (NLoS) component is physically due to many rays reflected or scattered from closely-spaced terrestrial obstacles within the vicinity of the BS, which are characterized by a very narrow beamwidth [9] and appear grouped or “clustered” in delay for sufficiently high elevation angles [54]. Henceforth, such rays are expected to have a very similar Doppler shift, as they all come from similar directions (narrow beamwidth), and a very similar delay (high elevation angle). Assuming that the UAV is moving (relative to the BS) at constant radial speed $v$ within the observation interval $[0,T_{0})$, with $T_{0}\triangleq N_{0}\,T_{\text{s }}\leq T_{\text{coh}}$, under the customary assumption that the communication bandwidth $B\sim 1/T_{\text{c}}$ is much smaller than the carrier frequency $f_{\text{carrier}}$, the far-field low-pass equivalent linear time-varying (LTV) time-delay impulse response of the UAV-to-BS channel can be expressed as [^6] Throughout the paper, the subscripts L and N indicate parameters referring to the LoS and NLoS components of the AU, respectively. (see, e.g., [53])
@@ 2.4 Receive signal model before CP removal
$x_{\text{A,C}}^{[p]}[n]$ $=\sum_{\ell=n-1}^{n}\sum_{q=0}^{P-1}u_{\text{A}}^{[q]}[\ell]$
$\hskip 42.67912pt\cdot\alpha_{\text{A,C}}\left[(n-\ell)P+(p-q)-d_{\text{A,C}}\right]$ (11)
$x_{\text{T},k}^{[p]}[n]$ $=\sum_{\ell=n-1}^{n}\sum_{q=0}^{P-1}u_{\text{T}}^{[q]}[\ell]$
$\hskip 42.67912pt\cdot\alpha_{\text{T},k}\left[(n-\ell)P+(p-q)-d_{\text{T},k}\right]$ (12)